OpenStudy (anonymous):

I need help understanding a problem. It says multiply and simplify if possible. sqrt3(sqrt12+sqrt6) How do I solve this? Please help

OpenStudy (mertsj):

\[\sqrt{3}(\sqrt{12}+\sqrt{6})=\sqrt{36}+\sqrt{18}=6+\sqrt{9(2)}=6+3\sqrt{2}\]